A pharmist wishes to mis a solution that is 5% Minoxidil. She has on hand 80ml of a 2% solution and wishes to add some 7% solution to obtain the desired 5% solution. How much 7% solution should she add?
Answer by checkley77(12844)

.07x+.02*80=.05(80+x)
.07x+1.6=4+.05x
.07x-.05x=4-1.6
.02x=2.4
x=2.4/.02
x=120 ml of thee 7% solution is needed.
Proof:
.07*120+.02*80=.05(80+120)
8.4+1.6=.05*200
10=10